OpenStudy (anonymous):

A line has a rise of 6 and a slope of 1/20. What is the run?

jimthompson5910 (jim_thompson5910):

slope = rise/run 1/20 = rise/run 1/20 = 6/run 1/20 = 6/x solve for x

OpenStudy (anonymous):

120?

jimthompson5910 (jim_thompson5910):

good